OpenStudy (anonymous):

How do you clear the equations of fractions?? E.G. Solve the Equation: x/2=10-x/3 PLEASE HELP!!!

OpenStudy (anonymous):

mult both sides by 6

OpenStudy (mathmagician):

x/2+x/3=10 (3x+2x)/6=10 5x/6=10 x=10*6/5=12

OpenStudy (anonymous):

3.x=60-2.x->5.x=60->X=12

OpenStudy (anonymous):

@mathmagician can you explain how u did that?